### v1

@@ 2. $i$
Appendix D Mixed Integer Quadratic Programming (MIQP)
This section develops a Mixed Integer Quadratic Programming (MIQP) approach to compute the exact value of the MRCI. We use this exact formulation to benchmark the accuracy and scalability of the heuristic optimization algorithm developed in Section 5.
D.1 MIQP Optimization
@@
D.2 Benchmarking: MIQP vs. Heuristic Approach
We compare the exact MIQP solution against the alternating maximization heuristic used in the main text (Section 5).
@@
The performance of the heuristic is near-exact. The mean absolute difference across all 2,058 tasks is $9.8\times 10^{-4}$ (approximately 0.001). Furthermore, 96.44% of all estimates fall within a strict 1% tolerance of the true global optimum. The maximum error observed in the entire benchmark was 0.085. These results provide strong evidence that the heuristic algorithm developed in Section 5 provides a high-quality approximation of the true MRCI while remaining computationally efficient for large datasets.
@@
Appendix E Additional Tables

### v2

@@ 2. $i$
Appendix E Mixed Integer Quadratic Programming (MIQP)
This section develops a Mixed Integer Quadratic Programming (MIQP) approach to compute the exact value of the MRCI. We use this exact formulation to benchmark the accuracy and scalability of the heuristic optimization algorithm developed in Section C.
E.1 MIQP Optimization
@@
E.2 Benchmarking: MIQP vs. Heuristic Approach
We compare the exact MIQP solution against the alternating maximization heuristic used in the main text (Section C).
@@
The performance of the heuristic is near-exact. The mean absolute difference across all 2,058 tasks is $9.8\times 10^{-4}$ (approximately 0.001). Furthermore, 96.44% of all estimates fall within a strict 1% tolerance of the true global optimum. The maximum error observed in the entire benchmark was 0.085. These results provide strong evidence that the heuristic algorithm developed in Section C provides a high-quality approximation of the true MRCI while remaining computationally efficient for large datasets.
@@
Appendix F Additional Tables